The following information is available for Department C for the month of August:

Units Cost
Work in process, August 1 (70% complete) 10,000
Direct materials $36,000
Direct labor 18,000
Manufacturing overhead 24,000
Total work in process, August 1 $78,000
Started in production during August 40,000
Costs added:
Direct materials $108,000
Direct labor 48,000
Manufacturing overhead     61,040
Total costs added during August $217,040
Work in process, August 31 (80% complete) 4,000
Units completed and transferred out 42,000


Materials are added at the beginning of the process. Inspection takes place at the 50 percent point in the conversion process to determine spoiled units. Round unit costs to two decimal places. The company uses the weighted average cost flow assumption.

Department C's cost per unit for materials using the weighted average method is

a.$3.60.

b.$2.88.

c.$3.43.

d.$2.16.
Answer #1

Equivalent unit of material = 10000+40000 = 50000 Unit

Cost per equivalent unit of material = (36000+108000)/50000 = 2.88

So answer is b) $2.88